### v1

@@ 4.4. Ablations and Model Analysis
Figure 8. Hyperparameter analysis of silhouette weight $\alpha$ and color weight $\beta$. Increasing $\alpha$ improves structural alignment up to a point, while larger $\beta$ enhances color fidelity.
5. Conclusion
@@
We propose IMAGGarment-1, a unified conditional diffusion framework for fine-grained garment generation with precise control over silhouette, color, and logo placement. By introducing mixed attention, color adapter, and the $A^{3}$ module, our framework explicitly disentangles global structure (silhouette and color) from local attributes (logo content and spatial placemen), enabling accurate spatial control and high-quality synthesis. To support this task, we construct GarmentBench, a large-scale benchmark with over 180K samples annotated with multi-level design conditions. Comprehensive experiments on both seen and unseen garments demonstrate that IMAGGarment-1 achieves state-of-the-art results in structure fidelity, color consistency, and logo controllability.
Supplementary Material
This supplementary material provides additional details and discussions to support the findings in the main paper. Section A introduces key notations and symbol definitions used throughout the paper. Section B explains the evaluation metrics, including their motivations and computational formulas. Section C presents an ethics statement discussing the potential risks and responsible use of our method. Section D includes extended results, such as more controllability visualization and more qualitative.
Appendix A Some Notations and Definitions
Table 3. Some notations and definitions.
Notation Definition
$\boldsymbol{t}$ Timestep
$\boldsymbol{Z_{t}}$ Latent feature at $t$ step
$\boldsymbol{Z_{m}}$ Output of mixed attention
$\boldsymbol{x_{0}}$ Real image
$\boldsymbol{x_{t}}$ Noisy data at $t$ step
$\boldsymbol{G}$ Garment image
$\boldsymbol{L}$ Logo image
$\boldsymbol{M}$ Mask image
$\boldsymbol{C_{g}}$ Feature of garment image
$\boldsymbol{C_{l}}$ Feature of logo image
$\boldsymbol{C_{m}}$ Feature of mask image
$\boldsymbol{C_{s}}$ Feature of silhouette image
$\boldsymbol{C_{c}}$ Feature of color image
$\boldsymbol{C_{t}}$ Feature of text prompt
$\boldsymbol{\tilde{I}_{c}}$ Generated garment image
$\boldsymbol{\theta}$ Diffusion model
$\boldsymbol{\epsilon}$ Gaussian noise
$\boldsymbol{\bar{\alpha}_{t}}$ Cumulative product of noise weights
$\boldsymbol{w}$ Guidance scale
$\boldsymbol{\alpha}$ Silhouette scale
$\boldsymbol{\beta}$ Color scale
Appendix B Evaluation Metrics
We employ five quantitative metrics to evaluate the performance of our method with respect to visual quality, structural and semantic consistency, and controllability. All metrics are computed between the generated image $I_{\text{gen}}$ and its corresponding ground truth $I_{\text{gt}}$, given a shared input condition set.
Fréchet Inception Distance (FID) (Heusel et al., 2017) measures the distribution-level similarity between generated and real images in the feature space of an Inception-V3 network. Let $(\mu_{r},\Sigma_{r})$ and $(\mu_{g},\Sigma_{g})$ denote the means and covariances of the real and generated image features, respectively. The metric is calculated as:
(12) $\text{FID}=\left\|\mu_{r}-\mu_{g}\right\|_{2}^{2}+\mathrm{Tr}\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right).$
Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) evaluates perceptual similarity between $I_{\text{gen}}$ and $I_{\text{gt}}$ using deep features from a pretrained network (AlexNet). Let $f^{l}(I)$ be the activation at layer $l$ and location $(h,w)$, and $w^{l}$ be a learned per-layer weight. LPIPS is computed as:
(13) $\text{LPIPS}(I_{\text{gen}},I_{\text{gt}})=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\left\|w^{l}\odot(f^{l}_{h,w}(I_{\text{gen}})-f^{l}_{h,w}(I_{\text{gt}}))\right\|_{2}^{2}.$
Color Structure Similarity (CSS) (Zeng et al., 2014) measures the distance between the generated and reference color palettes. Given palettes $P_{\text{gen}}$ and $P_{\text{gt}}$ with $K$ dominant colors extracted via K-means clustering, CSS is defined as:
(14) $\text{CSS}=\frac{1}{K}\sum_{i=1}^{K}\left\|P_{\text{gen}}^{(i)}-P_{\text{gt}}^{(i)}\right\|_{2}.$
Logo Location Accuracy (LLA) (Fujitake, 2024) quantifies the spatial deviation between the predicted and ground truth logo positions. Let $Rec_{gen}$ and $Rec_{gt}$ denote the rectangular areas of the logo in $I_{\text{gen}}$ and $I_{\text{gt}}$, respectively. The metric is computed as:
(15) $\text{LLA}=\frac{Intersection(Rec_{gen},Rec_{gt})}{Union(Rec_{gen},Rec_{gt})}.$
CLIPScore (Hessel et al., 2021) evaluates semantic alignment between the generated image and the input text prompt $T$. Let $f_{\text{img}}(\cdot)$ and $f_{\text{text}}(\cdot)$ be CLIP encoders for images and text. The metric is defined as the cosine similarity:
(16) $\text{CLIPScore}=\cos\left(f_{\text{img}}(I_{\text{gen}}),f_{\text{text}}(T)\right).$
Together, these metrics enable a rigorous and holistic assessment of fine-grained garment generation, covering realism, structure, color fidelity, spatial precision, and semantic consistency.
Appendix C Ethics Statement
This work introduces IMAGGarment-1, a controllable multi-stage diffusion framework for fine-grained garment image generation. While our method is developed to support applications such as personalized fashion design and virtual try-ons, we acknowledge potential ethical risks. In particular, the ability to manipulate real person images could be misused to generate deceptive or unauthorized content. Many human-centric image synthesis techniques share such concerns. To mitigate misuse, we advocate for responsible research and deployment practices. Our method is intended for academic and design-oriented purposes, and we encourage the use of detection, attribution, and content moderation tools alongside generative models. We believe that IMAGGarment-1’s contribution advances the field in a meaningful and controllable direction, and we support community efforts to ensure its safe and ethical use.
Appendix D Additional Results
More Controllability Visualization. Figure 9 provides additional visualizations of IMAGGarment-1’s controllability. These examples include challenging garment types such as textured tops, denim jeans, and skirts, further validating the model’s ability to achieve fine-grained and decoupled control over diverse attributes.
More Qualitative Results for IMAGGarment-1. Figure 10 displays additional results synthesized by IMAGGarment-1, demonstrating its robustness and high-quality garment generation across diverse test cases.
Figure 9. Controllability visualization of IMAGGarment-1. Each block varies one input condition while keeping others fixed. Top: Silhouette changes lead to consistent structural adaptation. Bottom: Color palette variation results in accurate color transfer. Examples include complex textures, denim, and skirts, demonstrating robustness across diverse garment types.
Figure 10. Additional qualitative results of IMAGGarment-1, demonstrating its robustness and visual fidelity across diverse garment styles and scenarios.

### v2

@@ IV-D Ablation Study
Figure 8: Hyperparameter analysis of silhouette weight $\alpha$ and color weight $\beta$.
IV-E More Results and Analysis
Controllability Analysis. We assess controllability by varying a single condition at a time (silhouette, color palette, or logo position) while keeping the others fixed. As shown in Fig. 7, each three column block visualizes the model’s response to one condition. Changing the silhouette (left block) yields garments that match the target shapes, indicating that the mixed attention module preserves structural alignment. Varying the color palette (middle block) produces the intended color distributions, validating the color adapter for color faithful generation. Adjusting the logo position (right block) achieves precise spatial relocation, showing that the $A^{3}$ module effectively injects spatial priors for local control. Overall, IMAGGarment provides fine-grained and decoupled control of garment attributes suitable for practical design workflows. Non-varied attributes remain stable across manipulations, reflecting minimal cross-condition interference and consistent editing behavior. Sequential composition of edits across attributes produces similar outcomes regardless of edit order, which suggests low inter-attribute coupling. Control fidelity also holds under moderate changes of viewpoint and background, supporting robustness in real design scenarios.
Hyperparameter Analysis. We study the effect of two key hyperparameters in Eq.1 and Eq.3: the silhouette guidance weight $\alpha$ and the color conditioning weight $\beta$. From Fig. 8, varying $\alpha$ directly impacts the model’s ability to follow the reference silhouette. When $\alpha$ is too low, the generated structure becomes blurry or deviates from the target shape; when too high, it may suppress color and text guidance. We empirically set $\alpha=0.6$ for balanced structural alignment. Similarly, the color weight $\beta$ controls the influence of the color palette. As $\beta$ increases, color consistency improves steadily, with $\beta=1.0$ yielding the best visual fidelity. Joint sweeps over $(\alpha,\beta)$ indicate a broad stability region around $\alpha\in[0.5,0.7]$ and $\beta\in[0.8,1.1]$, showing robustness to moderate mistuning. Interaction effects are mild: very large $\alpha$ slightly narrows the effective range of $\beta$, while very large $\beta$ can oversaturate colors and reduce shading nuance. We therefore adopt $\alpha=0.6$ and $\beta=1.0$ throughout all experiments.
V Conclusion
@@
We propose IMAGGarment, a unified conditional diffusion framework for fine-grained garment generation with precise control over silhouette, color, and logo placement. By introducing mixed attention, color adapter, and the $A^{3}$ module, our framework explicitly disentangles global structure (silhouette and color) from local attributes (logo content and spatial placement), enabling accurate spatial control and high-quality synthesis. To support this task, we construct GarmentBench, a large-scale benchmark with over 180K samples annotated with multi-level design conditions. Comprehensive experiments on both seen and unseen garments demonstrate that IMAGGarment achieves state-of-the-art results in structure fidelity, color consistency, and logo controllability. Code, models, and datasets are publicly available at https://github.com/muzishen/IMAGGarment.